OpenStudy (anonymous):

Lisa will make punch that is 25% fruit juice by adding pure fruit juice to a 2-liter mixture that is 10% pure fruit juice. How many liters of pure fruit juice does she need to add?

OpenStudy (lgbasallote):

hmm correct me if im wrong but the equation would be.. x liters of 100% fruit juice + 2 L of 10 % fruit juice = (x+2) liters of 25% fruit juice yes? that is how i remember it.. so... 1(x) + .10(2) = .25(x+2) lez multiply everything by 100 so no decimals... 100x + 20 = 25(x+2) 100x + 20 = 25x + 50 100x - 25x = 50-20 75x = 30 x = 30/75 = 6/15 = 2/5 2/5 = 0.4 so.. 0.4 Liters yes?

OpenStudy (anonymous):

Thanks!

OpenStudy (lgbasallote):

hmm so i guess that means that my answer was in the choices huh? yay i still remember this!